OpenStudy (anonymous):

Find any critcal numbers of the function: h(x)=sin^2x+cosx 0

OpenStudy (anonymous):

take derivative and set it equal to zero

OpenStudy (anonymous):

you are gonna have to use the chain rule with this problem

OpenStudy (anonymous):

or maybe rewrite using double angel therm

OpenStudy (anonymous):

The critical points of a function are the points where the derivative is zero and the points where the function is not differentiable. Since this function is differentiable everywhere in its domain (if the domain was\[[0, 2\pi],\]a closed interval, you would have to check the end points of the interval too) the only points that matter are where the derivative is zero. \begin{eqnarray*}&&h'(x) = 2\sin(x)\cos(x)-\sin(x) = \sin(x)(2\cos(x) - 1) = 0 \\ \Rightarrow&& \sin(x) = 0 \lor \cos(x) = \frac{1}{2} \Rightarrow x = \pi \lor x = \frac{\pi}{3} \lor x = \frac{5\pi}{3}.\end{eqnarray*}